Water moves through a pipe at 6 m/s under an absolute pressure of 330 kPa. The pipe narrows to one half of its original diameWater moves through a pipe at 6 m/s under an absolute pressure of 330 kPa. The pipe narrows to one half of its original diameter. What is the absolute pressure in the narrow part of the pipe?
Answer #1

Bernaullis formula - Po U² C P2 t +2₂ © V,? +2= + 7 zg २ zg P: pressure Ni unit weight water P = 330 kPa. V, = 6 m/s 2,= 22 oSo ให equation -ion 0 8 41 200x16 9:8IX ( - + + 24 8 2x9 :31 9:/x/ 2x9.8) Pa = 60 kf A หM